OpenStudy (anonymous):

what is the value of the expression -2(3x-2)+x+9 when x=-3

OpenStudy (anonymous):

-2(3(-3)-2)-3+9= -2(-9-2)-3+9= -2(-11)+6= 22+6=28

jimthompson5910 (jim_thompson5910):

Replace x with -3 and evaluate -2(3x-2)+x+9 -2(3(-3)-2)+(-3)+9 -2(-9-2)-3+9 -2(-11)-3+9 22-3+9 28 So -2(3x-2)+x+9 = 28 when x = -3